### v1

@@ Abstract
The fusion of a low-resolution hyperspectral image (LR-HSI) with a high-resolution multispectral image (HR-MSI) has emerged as an effective technique for achieving HSI super-resolution (SR). Previous studies have mainly concentrated on estimating the posterior distribution of the latent high-resolution hyperspectral image (HR-HSI), leveraging an appropriate image prior and likelihood computed from the discrepancy between the latent HSI and observed images. Low rankness stands out for preserving latent HSI characteristics through matrix factorization among the various priors. However, this method only enhances resolution within the dimensions of the two modalities. To overcome this limitation, we propose a novel continuous low-rank factorization (CLoRF) by integrating two neural representations into the matrix factorization, capturing spatial and spectral information, respectively. This approach enables us to harness both the low rankness from the matrix factorization and the continuity from neural representation in a self-supervised manner. Theoretically, we prove the low-rank property and Lipschitz continuity in the proposed continuous low-rank factorization. Experimentally, our method significantly surpasses existing techniques and achieves user-desired resolutions without the need for neural network retraining.
@@ 1 Introduction
Figure 1: The pipeline of CLoRF for arbitrary resolution. (A) Train the CLoRF. (B) Use the trained CLoRF to infer arbitrary resolutions of HSIs with given spatial and spectral coordinates. (C) An example of CLoRF for super-resolution on the Pavia University ($336\times 336\times 93$). The CLoRF is trained given its LR-MSI ($168\times 168\times 4$) and HR-HSI ($42\times 42\times 50$), then infers the original resolution. Bicubic interpolation directly upsamples from LR. (D) Visualize the spectrum of a random pixel from the results on (C).
Hyperspectral images (HSIs) have widespread applications across various fields due to their rich spectral information. The abundant spectral details offered by HSIs facilitate accurate scene interpretation and enhance the efficacy of numerous applications, including object classification \@BBOPcite\@BAP\@BBN[13\@BBCP and anomaly detection \@BBOPcite\@BAP\@BBN[15\@BBCP. However, the inherent trade-off between spectral and spatial resolution in HSI systems, constrained by hardware limitations, often results in HSIs with lower spatial resolution than RGB, panchromatic (PAN), and multispectral images (MSI). To enhance the spatial resolution of HSIs, a natural approach is to fuse LR-HSI and HR-MSI, known as hyperspectral and multispectral image fusion (HSI-MSI fusion). HSI-MSI fusion resembles MSI pansharpening, where low spatial resolution MSI is merged with high-resolution PAN imagery. However, directly applying these pansharpening methods to fuse HSI and MSI images suffers from challenges, as PAN images have limited spectral information, leading to spectral distortion \@BBOPcite\@BAP\@BBN[21\@BBCP. Consequently, numerous approaches tailored for HSI-MSI fusion are introduced, which can be generally categorized into model-based methods and deep learning-based models.
@@
Model-based methods leverage the low-rank structure of HSIs by characterizing their low-rankness through matrix factorization, decomposing the HSI matrix into the basis and coefficients or endmembers and abundances. Therefore, the principles of matrix factorization-based methods rely on appropriate prior information and the likelihood determined by the relationships between the latent HSI and the observed LR-HSI and HR-MSI. Given the known degradation model, most existing methods focus on modeling prior information for HSIs, including explicit and implicit methods. The explicit methods employ handcrafted explicit prior information, such as low-rankness \@BBOPcite\@BAP\@BBN[45, 34\@BBCP, smoothness \@BBOPcite\@BAP\@BBN[28\@BBCP, sparsity \@BBOPcite\@BAP\@BBN[11\@BBCP, and non-local similarity \@BBOPcite\@BAP\@BBN[9\@BBCP, to depict the prior distribution of the latent HSI for super-resolution (SR). Implicit smooth regularizations introduce basis functions to parameterize the prior information, extending the modeling to functional representations. For example, ? utilize non-negative matrix factorization parameterized by basis functions to reveal implicit smoothness. Furthermore, with the success of deep learning, ? implicitly introduce a denoising operator into the optimization framework to learn the deep image prior, which is shared across all HSIs using deep neural networks with different structures. Although these methods have achieved significant success in HSI-MSI fusion, manually designed explicit prior information may make the regularization terms problematic, resulting in significant computational complexity. Meanwhile, implicitly designed prior information sometimes becomes inappropriate for capturing the complex and detailed structures of HSIs. Moreover, these methods are restricted to fixed resolution and cannot be instantly applied to recovering HSIs with arbitrary resolution.
@@
Most deep learning-based methods typically involve supervised training on large datasets consisting of observed and ground truth data to learn a complex function mapping. Some representative deep learning methods are based on convolution neural network (CNN) for HSI-MSI fusion, such as the two-stream fusion network designed for HR-MSI and LR-HSI cite ?, [38, 18, 17, 36. With the remarkable performance of implicit neural representations (INRs) in continuous multi-dimensional data representation, some prior research cite ?, [37, 16 combined INR and CNN for HSI-MSI fusion. Although these methods have demonstrated promising results in fusion, they heavily rely on high-quality training data for supervised learning. Collecting high-quality ground truth data is extremely time-consuming and costly. In addition, unsupervised fusion methods cite ?, [46, 24, 40 are introduced, eliminating the need for expensive training data. However, these methods typically rely on complex network structures. Besides, they are also confined to fixed resolution, and inferring results at the user-desired resolutions requires retraining.
@@
To address the aforementioned issues, we propose an unsupervised HSI-MSI fusion framework for continuous low-rank factorization (CLoRF). Specifically, CLoRF integrates two implicit neural representations into the low-rank factorization, capturing continuous spatial and spectral information of HR-MSI and LR-HSI, respectively. Unlike classic discrete matrix factorization, we define the continuous matrix function factorization following tensor function cite ?, [22, which characterizes low-rankness in continuous representation. Each continuous function is a realization of INR parameterized by multi-layer perceptrons (MLPs). As MLPs are Lipschitz smooth, further characterizing smoothness in continuous representation. Compared to previous fusion methods, as shown in Fig. 1, CLoRF can achieve arbitrary resolution in both spatial and spectral domains without additional image information and retraining.
@@
    We experimentally demonstrate that CLoRF significantly surpasses existing techniques, which confirms its wide applicability and superiority.
  4. 4.
    We further expand the application of CLoRF to HSI and PAN image fusion (HSI-PAN fusion).
@@ 2.1 Implicit Neural Representation
INR offers a novel approach to representing implicitly defined, continuous, differentiable signals parameterized by neural networks cite ?, [29. INR has demonstrated remarkable performance in representing complex data structures, such as 3D reconstruction cite ?, [23, 29, 31 and 2D image super-resolution cite ?, [12, 3 and generation cite ?, [41, 5, etc. Recently, INR-based approaches have been explored for HSIs, such as HSI SR cite ?, [44, unmixing cite ?, [35, and fusion cite ?, [37, 8. Despite these commendable efforts, INR still encounters challenges in HSIs. For instance, HSIs consist of numerous spectral bands obtained through continuous imaging within a specific spectral range. However, INR itself may not possess sufficient stability to directly learn a valid continuous representation from the spectral domain of HSIs. In the fusion task, learning in both spatial and spectral domains is necessary; therefore, utilizing a single INR for learning representation may confront limited representation capabilities.
@@ 2.2 HSI-MSI Fusion via Continuous Representation
Several recent studies explore INRs for HSI-MSI fusion. For instance, ? proposes spatial-INR and spectral-INR for spatial and spectral resolution reconstruction, respectively. Besides, ? proposes an innovative fusion method that integrates CNN and INR. Based on them, ? develops two spectral-spatial INRs for arbitrary-resolution hyperspectral pansharpening. Although these fusion methods utilize spatial-spectral-based INRs, they significantly diverge from CLoRF. First, these methods rely on local implicit image functions cite ?, [5 for HSI SR. They employ a complex CNN network to encode spatial and spectral features, then feed these features and the three-dimensional coordinates of HSIs into an MLP to recover HSIs. Second, these HSI-MSI fusion methods are supervised learning, which heavily relies on pairs of images for training. Third, as they do not leverage the low-rank and smooth physical information inherent in HSIs, the computational cost of these INRs is intensive, primarily due to the large size of HSIs. Conversely, our method is model-based and unsupervised, finely encoding the low-rankness and smoothness into the continuous spatial-spectral factorization function. Consequently, this improves the stability of the continuous representation in HSIs and significantly reduces the complexity of the network structure and the computational cost. Therefore, our method is more feasible and general in various HSIs.
@@ 3.1 Problem Formulation
As HSIs generally have a low-rank structure, thus they lie in a low-dimensional subspace cite ?, [28, 47. The low-rank factorization aims to approximate a target matrix $\mathbf{Z}$ as a product of two matrices:
@@
where $\mathbf{E}\in\mathbb{R}^{L\times K}$ is a spectral dictionary and $\mathbf{A}\in\mathbb{R}^{K\times N}$ is a coefficient matrix, respectively. And $K\ll L$ represents a hyperparameter controlling the number of spectral bases. The low-rank factorization representation offers three main advantages. First, it maximizes the utilization of strong correlations among the spectral bands. Second, by keeping $K$ small (where $K\ll L$), the size of the spectral mode is reduced, thereby enhancing computational efficiency. Third, each row of matrix $\mathbf{A}$ can be linearly represented by the rows (bands) of matrix $\mathbf{Z}$, and the rows of matrix $\mathbf{A}$ maintain the spatial structures of matrix $\mathbf{Z}$.
@@ 3.2 Deep Continuous Low-rank Factorization Model
INR is widely adopted for learning continuous data representation, such as HSIs cite ?, [37, 8. However, simply utilizing a single INR to represent the HSI volume results in low efficiency and expensive computation, as it neglects the distinct low-rank structure of HSIs. Conversely, we propose a continuous low-rank factorization (CLoRF) model for learning HSI representation continuously and effectively. Our method fully explores the low-rank structure of HSIs by simultaneously learning its low-rank continuous representation by spatial and spectral INR. As a result, our approach effectively captures the low-rankness and smoothness of HSIs while overcoming the computational burden associated with existing INR-based methods cite ?, [37, 8 in HSIs.
As illustrated in Fig. 2, we present an overview of the proposed CLoRF. Specifically, CLoRF consists of two INRs for learning the low-rank continuous representation of HR-HSI. Inspired by the advance of Sinusoidal Representation Networks (SIRENs) cite ?, [29, we employ two SIRENs to estimate $\mathbf{E}$ and $\mathbf{A}$ in Eq.(3), respectively. Specifically, we denote one SIREN $\Psi_{\theta}(\cdot)$ parameterize by $\theta$ for approximating $\mathbf{E}$, and another SIREN $\Phi_{\alpha}(\cdot)$ parameterized by $\alpha$ for approximating $\mathbf{A}$, which is defined as follows:
@@ 3.3 Theoretical Analysis
In this section, we theoretically demonstrate that the low-rank and smooth regularizations are implicitly unified in continuous low-rank matrix factorization. Our analysis is inspired by the concept of tensor function factorization in cite ?, [22. Here, we start with rank factorization in the matrix computation field.
Theorem 1 (rank factorization cite ?, [25).
@@
Smoothness is another prevalent attribute in HSIs, such as the spatial and spectral smoothness of HSIs cite ?, [30. Here, we theoretically validate that our method incorporates implicit smooth regularization derived from the specific structures of MLPs.
@@
  • $\|\mathbf{X}\|_{1}$ is bounded by a positive constant $\eta_{1}$;
  • $\|\mathbf{W}_{i}^{1}\|_{1},\|\mathbf{W}_{i}^{2}\|_{1}$ are bounded by a positive constant $\eta_{2}$ for all $i$.
@@
where $\delta=\eta^{2d+1}\kappa^{2d-2}\zeta$, and $\eta=\max\{\eta_{1},\eta_{2}\}$.
The proof is shown in Appendix C. This smoothness is implicitly encoded with mild assumptions regarding nonlinear activation functions and weight matrices, which are readily attainable in a real-world implementation. For instance, commonly used activation functions such as ReLU, LeakyReLU, Sine, and Tanh are Lipschitz continuous. Additionally, cite ?, [22 effectively governs the level of smoothness by managing the structures of MLPs.

### v2

@@ Abstract
The fusion of a low-resolution hyperspectral image (LR-HSI) with a high-resolution multispectral image (HR-MSI) has emerged as an effective technique for achieving HSI super-resolution (SR). Previous studies have mainly concentrated on estimating the posterior distribution of the latent high-resolution hyperspectral image (HR-HSI), leveraging an appropriate image prior and likelihood computed from the discrepancy between the latent HSI and observed images. Low rankness stands out for preserving latent HSI characteristics through matrix factorization among the various priors. However, the primary limitation in previous studies lies in the generalization of a fusion model with fixed resolution scales, which necessitates retraining whenever output resolutions are changed. To overcome this limitation, we propose a novel continuous low-rank factorization (CLoRF) by integrating two neural representations into the matrix factorization, capturing spatial and spectral information, respectively. This approach enables us to harness both the low rankness from the matrix factorization and the continuity from neural representation in a self-supervised manner. Theoretically, we prove the low-rank property and Lipschitz continuity in the proposed continuous low-rank factorization. Experimentally, our method significantly surpasses existing techniques and achieves user-desired resolutions without the need for neural network retraining. Code is available at https://github.com/wangting1907/CLoRF-Fusion.
@@ 1 Introduction
Figure 1: The pipeline of CLoRF for arbitrary resolution. (A) Train the CLoRF. (B) Use the trained CLoRF to infer arbitrary resolutions of HSIs with given spatial and spectral coordinates. (C) An example of CLoRF for super-resolution on the Pavia University ($336\times 336\times 93$). The CLoRF is trained given its LR-MSI ($168\times 168\times 4$) and HR-HSI ($42\times 42\times 50$), then infers the original resolution. Bicubic interpolation directly upsamples from LR. In the spectral domain, each band has distinct brightness values, and bicubic interpolation estimates the missing bands by referencing adjacent ones. As a result, this can cause discrepancies in the color representation of the interpolated image, such as in band 60, when compared to the GT image. (D) Visualize the spectrum of a random pixel from the results on (C).
Hyperspectral images (HSIs) have widespread applications across various fields due to their rich spectral information. The abundant spectral details offered by HSIs facilitate accurate scene interpretation and enhance the efficacy of numerous applications, including object classification \@BBOPcite\@BAP\@BBN[16\@BBCP and anomaly detection \@BBOPcite\@BAP\@BBN[18\@BBCP. However, the inherent trade-off between spectral and spatial resolution in HSI systems, constrained by hardware limitations, often results in HSIs with lower spatial resolution than RGB, panchromatic (PAN), and multispectral images (MSI). To enhance the spatial resolution of HSIs, a natural approach is to fuse LR-HSI and HR-MSI, known as hyperspectral and multispectral image fusion (HSI-MSI fusion). HSI-MSI fusion resembles MSI pansharpening, where low spatial resolution MSI is merged with high-resolution PAN imagery. However, directly applying these pansharpening methods to fuse HSI and MSI images suffers from challenges, as PAN images have limited spectral information, leading to spectral distortion \@BBOPcite\@BAP\@BBN[28\@BBCP. Consequently, numerous approaches tailored for HSI-MSI fusion are introduced, which can be generally categorized into model-based methods and deep learning-based models.
@@
Model-based methods leverage the low-rank structure of HSIs by characterizing their low-rankness through matrix factorization, decomposing the HSI matrix into the basis and coefficients or endmembers and abundances. Therefore, the principles of matrix factorization-based methods rely on appropriate prior information and the likelihood determined by the relationships between the latent HSI and the observed LR-HSI and HR-MSI. Given the known degradation model, most existing methods focus on modeling prior information for HSIs, including explicit and implicit methods. The explicit methods employ handcrafted explicit prior information, such as low-rankness \@BBOPcite\@BAP\@BBN[57, 42\@BBCP, smoothness \@BBOPcite\@BAP\@BBN[36\@BBCP, sparsity \@BBOPcite\@BAP\@BBN[14\@BBCP, and non-local similarity \@BBOPcite\@BAP\@BBN[11\@BBCP, to depict the prior distribution of the latent HSI for super-resolution (SR). Implicit smooth regularizations introduce basis functions to parameterize the prior information, extending the modeling to functional representations. For example, ? utilize non-negative matrix factorization parameterized by basis functions to reveal implicit smoothness.
Compared to matrix-based methods, tensor-based methods can directly process HSI data and have garnered significant attention cite ?, [12, 7. For example, a nonlocal sparse tensor factorization method based on Tucker decomposition was introduced in cite ?, [26, which factorizes an HSI into a sparse core tensor multiplied by dictionaries along both the spatial and spectral dimensions. In cite ?, [55, a novel fusion model is proposed within the tensor ring decomposition framework, rather than using subspace decomposition-based fidelity terms. Although these tensor decomposition-based methods are effective in preserving the spectral-spatial correlation of HSI, they face the challenge of dimensionality catastrophe. Furthermore, with the success of deep learning, ? implicitly introduce a denoising operator into the optimization framework to learn the deep image prior, which is shared across all HSIs using deep neural networks with different structures. Although these methods have achieved significant success in HSI-MSI fusion, manually designed explicit prior information may make the regularization terms problematic, resulting in significant computational complexity. Meanwhile, implicitly designed prior information sometimes becomes inappropriate for capturing the complex and detailed structures of HSIs. Moreover, these methods are restricted to fixed resolution and cannot be instantly applied to recovering HSIs with arbitrary resolution.
@@
Most deep learning-based methods typically involve supervised training on large datasets consisting of observed and ground truth data to learn a complex function mapping. Some representative deep learning methods are based on convolution neural network (CNN) for HSI-MSI fusion, such as the two-stream fusion network designed for HR-MSI and LR-HSI cite ?, [46, 23, 22, 44. With the remarkable performance of implicit neural representations (INRs) in continuous multi-dimensional data representation, some prior research cite ?, [45, 20 combined INR and CNN for HSI-MSI fusion. Although these methods have demonstrated promising results in fusion, they heavily rely on high-quality training data for supervised learning. Collecting high-quality ground-truth data is extremely time-consuming and costly. In addition, unsupervised fusion methods cite ?, [59, 31, 48 are introduced, eliminating the need for expensive training data. However, these methods typically rely on complex network structures. Besides, they are also confined to fixed resolution, and inferring results at the user-desired resolutions requires retraining. Arbitrary-resolution HSI-MSI fusion is an emerging area within the field of image fusion. Unlike traditional fusion tasks, which typically merge images at fixed resolutions, arbitrary-resolution fusion leverages both HR-MSI and LR-HSI to generate fused outputs at any desired spatial and spectral resolutions cite ?, [44, 21. This advanced fusion technique provides greater flexibility in adjusting both the spatial and spectral resolutions of the resulting HR-HSI, offering significant potential for a wide range of applications cite ?, [21, such as object detection cite ?, [33 and land use/cover classification cite ?, [16. However, arbitrary-resolution hyperspectral image fusion presents greater challenges compared to standard fusion methods.
@@
To address the aforementioned issues, we propose an unsupervised HSI-MSI fusion framework for continuous low-rank factorization (CLoRF). Specifically, CLoRF integrates two implicit neural representations into the low-rank factorization, capturing continuous spatial and spectral information of HR-MSI and LR-HSI, respectively. Unlike classic discrete matrix factorization, we define the continuous matrix function factorization following tensor function cite ?, [29, which characterizes low-rankness in continuous representation. Each continuous function is a realization of INR parameterized by multi-layer perceptrons (MLPs). As MLPs are Lipschitz smooth, further characterizing smoothness in continuous representation. Compared to previous fusion methods, as shown in Fig. 1, CLoRF can achieve arbitrary resolution in both spatial and spectral domains without additional image information and retraining.
@@
    We experimentally demonstrate that CLoRF significantly surpasses existing techniques, confirming its wide applicability and superiority and further expanding its application to HSI and PAN image fusion (HSI-PAN fusion).
@@ 2.1 Implicit Neural Representation
INR offers a novel approach to representing implicitly defined, continuous, differentiable signals parameterized by neural networks cite ?, [37. INR has demonstrated remarkable performance in representing complex data structures, such as 3D reconstruction cite ?, [30, 37, 39 and 2D image super-resolution cite ?, [15, 3 and generation cite ?, [49, 6, etc. Recently, INR-based approaches have been explored for HSIs, such as HSI SR cite ?, [56, unmixing cite ?, [43, and fusion cite ?, [45, 10. Despite these commendable efforts, INR still encounters challenges in HSIs. For instance, HSIs consist of numerous spectral bands obtained through continuous imaging within a specific spectral range. However, INR itself may not possess sufficient stability to directly learn a valid continuous representation from the spectral domain of HSIs. In the fusion task, learning in both spatial and spectral domains is necessary; therefore, utilizing a single INR for learning representation may confront limited representation capabilities.
@@ 2.2 HSI-MSI Fusion via Continuous Representation
Several recent studies explore INRs for HSI-MSI fusion. For instance, ? proposes spatial-INR and spectral-INR for spatial and spectral resolution reconstruction, respectively. Besides, ? proposes an innovative fusion method that integrates CNN and INR. Based on them, ? develops two spectral-spatial INRs for arbitrary-resolution hyperspectral pansharpening. Although these fusion methods utilize spatial-spectral-based INRs, they significantly diverge from CLoRF. First, these methods rely on local implicit image functions cite ?, [6 for HSI SR. They employ a complex CNN network to encode spatial and spectral features, then feed these features and the three-dimensional coordinates of HSIs into an MLP to recover HSIs. Second, these HSI-MSI fusion methods are supervised learning, which heavily relies on pairs of images for training. Third, they do not perform low-rank matrix decomposition on HR-HSI, nor do they leverage the low-rank and smooth physical information inherent in HSIs. As a result, the computational cost of these INRs is high, primarily due to the large size of HSIs, which results in large three-dimensional coordinates. Conversely, our method is model-based and unsupervised, finely encoding the low-rankness and smoothness into the continuous spatial-spectral factorization function. Consequently, this improves the stability of the continuous representation in HSIs and significantly reduces the complexity of the network structure and the computational cost. Therefore, our method is more feasible and general in various HSIs.
@@ 3.1 Problem Formulation
As HSIs generally have a low-rank structure, thus they lie in a low-dimensional subspace cite ?, [36, 60. The low-rank factorization aims to approximate a target matrix $\mathbf{Z}$ as a product of two matrices:
@@
where $\mathbf{E}\in\mathbb{R}^{L\times K}$ is a spectral dictionary and $\mathbf{A}\in\mathbb{R}^{K\times N}$ is a coefficient matrix, respectively. And $K\ll L$ represents a hyperparameter controlling the number of spectral bases. The low-rank factorization representation offers three main advantages. First, it maximizes the utilization of strong correlations among the spectral bands. Second, by keeping $K$ small (where $K\ll L$), the size of the spectral mode is reduced, thereby enhancing computational efficiency. Third, each column of matrix $\mathbf{Z}$ can be linearly represented by the columns of matrix $\mathbf{E}$ using the coefficients in matrix $\mathbf{A}$. The rows of matrix $\mathbf{A}$ maintain the spatial structures of matrix $\mathbf{Z}$. Note that (3) is not a unique factorization for $\mathbf{Z}$. One could obtain another pairs $\hat{\mathbf{E}}=\mathbf{E}\mathbf{B}$ and $\hat{\mathbf{A}}=\mathbf{B}^{-1}\mathbf{A}$, with any inverse matrix $\mathbf{B}\in\mathbb{R}^{K\times K}$.
@@ 3.2 Deep Continuous Low-rank Factorization Model
INR is widely adopted for learning continuous data representation, such as HSIs cite ?, [45, 10. However, simply utilizing a single INR to represent the HSI volume results in low efficiency and expensive computation, as it neglects the distinct low-rank structure of HSIs. Conversely, we propose a continuous low-rank factorization (CLoRF) model for learning HSI representation continuously and effectively. Our method fully explores the low-rank structure of HSIs by simultaneously learning its low-rank continuous representation by spatial and spectral INR. As a result, our approach effectively captures the low-rankness and smoothness of HSIs while overcoming the computational burden associated with existing INR-based methods cite ?, [45, 10 in HSIs.
As illustrated in Fig. 2, we present an overview of the proposed CLoRF. Specifically, CLoRF consists of two steps: low-rank decomposition and learning. The low-rank decomposition breaks down the HSI data space into two smaller subspaces: spatial basis $\mathbf{A}$ and spectral transformation $\mathbf{E}$. Additionally, the spatial and spectral components are parameterized by two neural networks, using two INRs to learn the low-rank continuous representation of the HR-HSI. Inspired by the advance of Sinusoidal Representation Networks (SIRENs) cite ?, [37, we employ two SIRENs to estimate $\mathbf{E}$ and $\mathbf{A}$ in Eq.(3), respectively. Specifically, we denote one SIREN $\Psi_{\theta}(\cdot)$ parameterize by $\theta$ for approximating $\mathbf{E}$, and another SIREN $\Phi_{\alpha}(\cdot)$ parameterized by $\alpha$ for approximating $\mathbf{A}$, which is defined as follows:
@@ 3.3 Theoretical Analysis
In this section, we theoretically demonstrate that the low-rank and smooth regularizations are implicitly unified in continuous low-rank matrix factorization. Our analysis is inspired by the concept of tensor function factorization in cite ?, [29. Here, we start with rank factorization in the matrix computation field.
Theorem 1 (rank factorization cite ?, [32).
@@
Remark 1.
1) In Theorem 1, the low-rank matrix decomposition definitely exists but is non-unique. This is because the representation of eigenvectors in SVD is not unique, which leads to different $\mathbf{U}$ and $\mathbf{U}^{\prime}$, as well as $\mathbf{V}$ and $\mathbf{V}^{\prime}$ in the decomposition, such as $\mathbf{U}^{\prime}=c\mathbf{U}$, $\mathbf{V}^{\prime}=\mathbf{V}/c$, where $c$ is a nonzero scalar.
2) CLoRF comprises two subnetworks with identical architectures: Spatial-INR and Spectral-INR. This network framework, equipped with low-rank and smooth priors, learns to generate a spectral dictionary and spatial coefficient matrix, enabling a low-rank representation of HSIs through network training. The solution spaces of spatial and spectral basis are constrained by the parameters of the INRs. Furthermore, by incorporating TV prior into the loss function, the solution space of the spatial coefficient matrix is further restricted, effectively reducing the ambiguity in the low-rank factorization.
Smoothness is another prevalent attribute in HSIs, such as the spatial and spectral smoothness of HSIs cite ?, [38. Here, we theoretically validate that our method incorporates implicit smooth regularization derived from the specific structures of MLPs.
@@
  • $\|\mathbf{W}_{i}^{1}\|_{1},\|\mathbf{W}_{i}^{2}\|_{1}$ are bounded by a positive constant $\eta$ for all $i$.
@@
where $\delta=\eta^{2d+1}\kappa^{2d-2}\zeta$, and $\zeta=\max\{\|\mathbf{s}_{1}\|_{1},|b_{1}|\}$.
The proof is shown in Appendix C. This smoothness is implicitly encoded with mild assumptions regarding nonlinear activation functions and weight matrices, which are readily attainable in a real-world implementation. For instance, we utilized the sine activation function, while also ensuring that the weights in the MLP network cite ?, [37 remain bounded.
Remark 2.
In Theorem 4, we observe that the degree of smoothness, denoted as $\delta$, is associated with the Lipschitz constant $\kappa$ and the upper bound $\eta$ of the weight matrices. Therefore, we can manipulate two variables in practice to achieve a balance in implicit smoothness:
1) We utilize the Sine function $\sigma(\cdot)=\sin(\omega_{0}\cdot)$ as the nonlinear activation function in the MLPs. Since the Sine function is Lipschitz continuous, we can effectively adjust its Lipschitz constant $\kappa$ by varying the value of $\omega_{0}$. Specifically, a smaller $\omega_{0}$ results in a lower Lipschitz constant $\kappa$, leading to smoother outcomes.
2) To manage the upper bound $\eta$ of the MLP weight matrices, we can adjust the trade-off parameter in the energy regularization of the MLP weights, commonly referred to as weight decay in contemporary deep learning optimizers. This approach allows us to control the strength of $\eta$.
Remark 3.
Assume the assumptions in Theorem 4 are satisfied. We define $f(\cdot):=[\Phi_{\alpha},\Psi_{\theta}](\cdot)$. Then, for any matrix $\mathbf{M}\in\mathcal{S}[f]$ that is sampled using coordinates vectors $\mathbf{s}\in\mathcal{A}_{f}$, $t\in\mathcal{Z}_{f}$, where $\mathcal{S}[f]$ represents the set of sampled matrices from the matrix function $f(\cdot)$ as defined in Definition 1, the following inequalities hold for $(i,j),i=1,2,\cdots,n_{1},j=1,2,\cdots n_{2}:$
$\small|\mathbf{M}_{(\mathbf{s}_{i},t_{j})}-\mathbf{M}_{(\mathbf{s}_{i-1},t_{j-1})}|\leq\delta\|\mathbf{s}_{i}-\mathbf{s}_{i-1}\|_{1}+\delta|t_{j}-t_{j-1}|,$ (8)
where $\delta=\eta^{2d+1}\kappa^{2d-2}\zeta$, and $\zeta=\max\{\|\mathbf{s}_{1}\|_{1},|b_{1}|\}$.
Remark 3 states that for any sampled matrix $\mathbf{M}\in S[f]$, the difference between adjacent elements is constrained by the distance between the corresponding coordinates, with the inclusion of a constant factor.